Using integrating factor, solve the differential equation frac {dy}{dx}+y=e^{x}
1
Expert's answer
2019-10-11T09:41:30-0400

The linear first order differential equation:


"\frac {dy}{dx} + P(x)y = Q(x)"


has the integrating factor (IF) : "IF= e^{\int P (x) dx}"


(1) "\frac {dy}{dx}+y=e^{x}" .


Integrating factor:


"P(x) = 1" .


Integrating factor, "IF= \int e^{\int P (x) dx} = e^{\int dx}= e^{x}"

"IF= e^{x}"


Multiply equation by IF:


"e^x\frac {dy}{dx} + e^xy = e^xe^x"

So,


"e^x\frac {dy}{dx} + e^xy = e^{2x}" ,


i.e. "\frac {d}{dx}[ e^xy] = e^{2x}"


Integrate:

"e^xy = \int e^{2x} dx"

i.e.

"e^xy = \frac {1}{2}e^{2x} + C ,"


"y = e^{-x}[\frac {1}{2}e^{2x} + C]."

.



Therefore, general solution of a Differential Equation (1):



"y = \frac {1}{2}e^{x} + Ce^{-x}"

.